OpenStudy (anonymous):

Review the work-energy theorem and apply it to a simple problem. If you push a particle of mass in the direction in which it is already moving, you expect the particle's speed to increase. If you push with a constant force , then the particle will accelerate with acceleration (from Newton's 2nd law). What is the final speed of the particle? Express your answer in terms of Kf and M.

OpenStudy (anonymous):

then the __velocity___ of the particle will increase by an amount at. then the __energy___ of the particle will increase by FD. Kf = Ki + W Kf = (1/2)*M*Vi^2 + F*D Vf^2 = 2*Kf/M Vf = sqrt(2*Kf/M)